The t statistic for a test of ?0:?=58 ??:?≠58 based on n = 6 observations has the value t = -1.93. (a) What are the degrees of freedom for this statistic? (b) Using the appropriate table in your formula packet, bound the p-value as closely as possible: ____< p-value<______
Answer #1

From the given information,

n= 6

t= -1.93

a. Degrees of freedom= n-1= 6-1 = 5

b. P-value= 0.11149

Hence, 0.11 < p-value < 0.12

Thank you.